OpenStudy (anonymous):

Solve the equation. 2.4=-13.6+(-3.4c)+1.7c

OpenStudy (anonymous):

A. –8 B. 2 C. –10 D. –4

OpenStudy (anonymous):

should be C...... you have different thing on the question?

OpenStudy (anonymous):

if 3.4=-13.6+(-3.4c)+1.7c 3.4+13.6=-3.4c+1.7c 17=-1.7c c=-10